# Leveraging Anthropometric Measurements to Improve Human Mesh Estimation and Ensure Consistent Body Shapes

### v3

@@ 9 Evaluating A2B Models
Figure 9: Visualization of the A2B evaluation and training procedures. The training part is highlighted with thicker arrows. During training, the $\beta$ parameter error is used. For evaluatations, the $\beta$ parameter error and the anthropometric error are calculated.
@@ 11 Generation of Pseudo GT Anthropometric Measurements
SMPLer-X 86.02 SMPLer-X 85.89 85.69 86.03 85.99 86.04
SMPLer-X FT 79.09 SMPLer-X FT 78.92 78.88 79.44 79.37 79.44
SMPLer-X FT - GT 65.63 65.84 64.71 64.76 -
SMPLer-X FT - SMPLer-X 73.41 73.29 73.65 73.63 73.66
IK-UU 67.54 IK-UU 66.92 66.60 67.25 67.12 67.16
IK-UU - SMPLer-X 63.80 63.64 63.81 63.78 63.82
IK-UU - SMPLer-X FT 69.46 69.27 69.70 69.63 69.69
IK-UU - GT 56.44 56.56 55.14 55.19 -
Table 10: MPJPE results in mm for the test split of ASPset. Results are given for different methods and replaced $beta$ parameters with A2B results (columns NN/SVR) or the median of the original $\beta$ parameters from the model noted in the measure column. SMPLer-X FT stands for the best finetuned variant of SMPLer-X (finetuned with the meshes obtained from IK executed on the GT 3D joints). The orig column contains the results without replaced $\beta$ parameters. We highlight the best result for each model and the best option for the combination of IK-UU pose and SMPLer-X $\beta$ parameters, since this combination outperforms the original IK-UU result, too.
@@
We consider this strategy as a valid method for evaluations, since our main goal is to improve the HME performace as much as possible with only marginal overhead. Our main focus is sports, which contains extreme poses that lets existing HME models fail, sometimes even to detect a human at all. Examples can be found in the supplementary videos. As professional athletes are measured anyway, the additional effort for the measurements is negligible in this context.
@@ 12 Inverse Kinematics
13 Finetuning HME Models with Pseudo GT Meshes
Finetuning existing HME models on pure 3D joints datasets is not possible, since they need mesh annotations for training. However, with IK, we can generate pseudo GT meshes. We exemplary test a finetuning of SMPLer-X on ASPset with this approach. Experiments show that using their finetuning script with 1.6M iterations leads to worse results than the results without finetuning. Therefore, we reduce the number of iterations with early stopping and achieve better results with finetuning only for 32K iterations.
The results shown in Table 10 prove that finetuning on IK generated meshes can lead to a significant improvement of the scores. Replacing the $\beta$ parameters of the finetuned results with the A2B $\beta$ parameters boosts the performance even more. These are the best results achieved with any existing HME model throughout this study.
Moreover, we experiment with using the SMPLer-X body shape parameters combined with the poses estimated by IK applied to the UU results (see last two rows of Table 10). Using the $\beta$ parameters from SMPLer-X leads to a slightly better result than the original 3D joint based result (without IK). This evaluation shows that 3D HPE models are better in precisely locating the joints of humans than HME models, but HME models are better in estimating the shape of humans. We also try to use the $\beta$ parameters of the finetuned variant together with the UU IK poses like before. However, this resulted in a performance drop compared to the body shape parameters from the original SMPLer-X without finetuning. These experiments show that finetuning HME models on pseudo ground truth leads to a better performance regarding the keypoints, but the estimated $\beta$ parameters have worse quality. This can further be proven by replacing the $\beta$ parameters from the finetuned SMPLer-X variant with the $\beta$ parameters from the not finetuned model, which results in a performance gain of over 5 mm compared to the original results from the finetuned version (rows 2 and 4 in Tab. 10). However, our method using the UU IK poses and the A2B body shape parameters with GT anthropometric measurments achieves the overall best results.
We provide a comprehensive summary and visualization of all results on the ASPset dataset in Section 14. This includes results of existing HME models, results of our approach, and the finetuning results.
@@ 14 Summary of the Results
We execute a multitude of experiments with different combinations of pose and shape parameters. Figure 10 summarizes the results with their pose and shape origins for ASPset. In general, the poses estimated by IK based on the UU results (red branch in Fig. 10) are more precise than the poses estimated by SMPLer-X (light blue branch in Fig. 10). Further, the body shape parameters from our A2B models with GT anthropometric measurements (green boxes in Fig. 10) achieve the best results for all poses. We provide more qualitative examples comparing SMPler-X with this approach in the supplementary video. Without access to the GT, all models benefit slightly from A2B model results with the median anthropometric measurements from B2A of the estimated meshes by the respective model (boxes with same color for all three branches in Fig. 10). Moreover, SMPLer-X A2B body shape parameters perform best when analyzing body shapes without GT access (light blue boxes in Fig. 10). Finetuning SMPLer-X with IK created meshes (dark blue branch in Fig. 10) improves the performance of SMPLer-X, although the quality of the body shape deteriorates. This can be seen as by comparing the shapes from SMPLer-X and finetuned SMPLer-X (dark blue and light blue boxes in Fig. 10) with finetuned and IK poses.
Since fit3D is a larger dataset, finetuning UU works better, which further leads to better IK meshes with an MPJPE of 37.02 mm. Enforcing consistent meshes with GT or IK A2B shape parameters decreases the performance slightly in this case. However, A2B shape parameters achieve slightly better scores than median values. This also holds for OSX and Multi-HMR. Overall, the approach with UU, IK, and A2B body shape parameters achieves an over 33 mm lower MPJPE than any HME model. The same also holds for the MVE, which can be improved by over 30 mm with our approach. The scores can be found in the main paper.
@@
Figure 11: Example frames from our supplementary videos. It shows qualitative results of our approach compared to MultiHMR for fit3d (left) and qualitative results of SMPLer-X and our approach for example frames from ASPSet (right). In fit3d visualizations, we display the GT meshes in green and the estimated meshes in grey. The GT joints are also displayed in green while the estimated joints from our model are visualized in blue blue. The MultiHMR joints are shown in red. Corresponding joints are connected. We display the exact MPJPE values in the top left of each frame. Recall that the visualization is in 2D, but the evaluation is in 3D. Therefore, sometimes the MPJPE values may seem odd. In the lower part, we show the estimated body shapes in T-pose. The GT body shape is shown in green and the estimated body shape from our model in blue. The MultiHMR body shape is shown in red. For ASPset visualizations, we display the estimated meshes and the GT and estimated joints. GT joints are shown in green, estimated joints from our model in blue. and the SMPLer-X joints in red. Corresponding joints are connected. In the lower part, we show the GT and estimated joints in the same way, but without the mesh and image to reduce distraction. We further display the MPJPE values.

### v4

@@ 9 Evaluating A2B Models
Figure 9: Visualization of the A2B evaluation and training procedures. The training part is highlighted with thicker arrows. During training, the $\beta$ parameter error is used. For evaluations, the $\beta$ parameter error and the anthropometric error are calculated.
In the main paper, we test our A2B models on the AGORA [26] dataset and randomly sampled body shapes. Since AGORA is a synthetic dataset, it might not reflect the real world. The same holds for randomly sampled body shapes. Therefore, we additionally test our best A2B models on the real-world SSP-3D dataset [33] which consists of diverse body shapes. We display the results in Table 10.
[…]
m f n m f n
NN 1.73 0.97 2.74 0.634 0.803 0.968
SVR 0.13 0.0039 0.066 0.167 0.114 0.182
Table 10: Results of our A2B models on the SSP-3D dataset using n(eutral), m(ale) and f(emale) meshes.
All A2B models accurately estimate the diverse real-world body shapes with low error.
@@ 11 Generation of Pseudo GT Anthropometric Measurements
SMPLer-X 86.02 SMPLer-X 85.89 85.69 86.02 85.99 86.04
SMPLer-X FT 79.09 SMPLer-X FT 78.92 78.88 79.59 79.37 79.44
SMPLer-X FT - GT 65.63 65.84 64.77 64.76 -
SMPLer-X FT - SMPLer-X 73.41 73.29 73.79 73.63 73.66
IK-UU 67.54 IK-UU 66.92 66.60 67.28 67.12 67.16
IK-UU - SMPLer-X 63.80 63.64 63.92 63.78 63.82
IK-UU - SMPLer-X FT 69.46 69.27 69.83 69.63 69.69
IK-UU - GT 56.44 56.56 55.18 55.19 -
Table 11: MPJPE results in mm for the test split of ASPset. Results are given for different methods and replaced $beta$ parameters with A2B results (columns NN/SVR) or the median of the original $\beta$ parameters from the model noted in the measure column. SMPLer-X FT stands for the best fine-tuned variant of SMPLer-X (fine-tuned with the meshes obtained from IK executed on the GT 3D joints). The orig column contains the results without replaced $\beta$ parameters. We highlight the best result for each model and the best option for the combination of IK-UU pose and SMPLer-X $\beta$ parameters, since this combination outperforms the original IK-UU result, too.
@@
We consider this strategy as a valid method for evaluations, since our main goal is to improve the HME performance as much as possible with only marginal overhead. Our main focus is sports, which contains extreme poses that let existing HME models fail, sometimes even to detect a human at all. Examples can be found in the supplementary videos. As professional athletes are measured anyway, the additional effort for the measurements is negligible in this context.
@@ 12 Inverse Kinematics
13 Fine-tuning HME Models with Pseudo GT Meshes
Fine-tuning existing HME models on pure 3D joints datasets is not possible, since they need mesh annotations for training. However, with IK, we can generate pseudo GT meshes. We exemplary test a fine-tuning of SMPLer-X on ASPset with this approach. Experiments show that using their fine-tuning script with 1.6M iterations leads to worse results than the results without fine-tuning. Therefore, we reduce the number of iterations with early stopping and achieve better results with fine-tuning only for 32K iterations.
The results shown in Table 11 prove that fine-tuning on IK generated meshes can lead to a significant improvement of the scores. Replacing the $\beta$ parameters of the fine-tuned results with the A2B $\beta$ parameters boosts the performance even more. These are the best results achieved with any existing HME model throughout this study.
Moreover, we experiment with using the SMPLer-X body shape parameters combined with the poses estimated by IK applied to the UU results (see last two rows of Table 11). Using the $\beta$ parameters from SMPLer-X leads to a slightly better result than the original 3D joint based result (without IK). This evaluation shows that 3D HPE models are better in precisely locating the joints of humans than HME models, but HME models are better in estimating the shape of humans. We also try to use the $\beta$ parameters of the fine-tuned variant together with the UU IK poses like before. However, this resulted in a performance drop compared to the body shape parameters from the original SMPLer-X without fine-tuning. These experiments show that fine-tuning HME models on pseudo ground truth leads to a better performance regarding the keypoints, but the estimated $\beta$ parameters have worse quality. This can further be proven by replacing the $\beta$ parameters from the fine-tuned SMPLer-X variant with the $\beta$ parameters from the not fine-tuned model, which results in a performance gain of over 5 mm compared to the original results from the fine-tuned version (rows 2 and 4 in Tab. 11). However, our method using the UU IK poses and the A2B body shape parameters with GT anthropometric measurements achieves the overall best results.
We provide a comprehensive summary and visualization of all results on the ASPset dataset in Section 14. This includes results of existing HME models, results of our approach, and the fine-tuning results.
@@ 14 Summary of the Results
Figure 11: Example frames from our supplementary videos. It shows qualitative results of our approach compared to MultiHMR for fit3d (left) and qualitative results of SMPLer-X and our approach for example frames from ASPset (right). In fit3d visualizations, we display the GT meshes in green and the estimated meshes in gray. The GT joints are also displayed in green while the estimated joints from our model are visualized in blue. The MultiHMR joints are shown in red. Corresponding joints are connected. We display the exact MPJPE values in the top left of each frame. Recall that the visualization is in 2D, but the evaluation is in 3D. Therefore, sometimes the MPJPE values may seem odd. In the lower part, we show the estimated body shapes in T-pose. The GT body shape is shown in green and the estimated body shape from our model in blue. The MultiHMR body shape is shown in red. For ASPset visualizations, we display the estimated meshes and the GT and estimated joints. GT joints are shown in green, estimated joints from our model in blue. and the SMPLer-X joints in red. Corresponding joints are connected. In the lower part, we show the GT and estimated joints in the same way, but without the mesh and image to reduce distraction. We further display the MPJPE values.
We execute a multitude of experiments with different combinations of pose and shape parameters. Figure 10 summarizes the results with their pose and shape origins for ASPset. In general, the poses estimated by IK based on the UU results (red branch in Fig. 10) are more precise than the poses estimated by SMPLer-X (light blue branch in Fig. 10). Further, the body shape parameters from our A2B models with GT anthropometric measurements (green boxes in Fig. 10) achieve the best results for all poses. We provide more qualitative examples comparing SMPLer-X with this approach in the supplementary video. Without access to the GT, all models benefit slightly from A2B model results with the median anthropometric measurements from B2A of the estimated meshes by the respective model (boxes with same color for all three branches in Fig. 10). Moreover, SMPLer-X A2B body shape parameters perform best when analyzing body shapes without GT access (light blue boxes in Fig. 10). Fine-tuning SMPLer-X with IK created meshes (dark blue branch in Fig. 10) improves the performance of SMPLer-X, although the quality of the body shape deteriorates. This can be seen as by comparing the shapes from SMPLer-X and fine-tuned SMPLer-X (dark blue and light blue boxes in Fig. 10) with fine-tuned and IK poses.
Since fit3D is a larger dataset, fine-tuning UU works better, which further leads to better IK meshes with an MPJPE of 37.02 mm. Enforcing consistent meshes with GT or IK A2B shape parameters decreases the performance slightly in this case. However, A2B shape parameters achieve slightly better scores than median values. This also holds for OSX and Multi-HMR. Overall, the approach with UU, IK, and A2B body shape parameters achieves an over 33 mm lower MPJPE than any HME model. The same also holds for the MVE, which can be improved by over 30 mm with our approach. The scores can be found in the main paper.